# Personalized Federated Learning under Model Dissimilarity Constraints

### v1

@@ 3.3 Algorithm
To solve the Karula problem (1), we need an appropriate algorithm. Here, we present a projected gradient algorithm that need only a small subset of clients to participate in each training round. To eliminate the error introduced by partial participation in each training round, our algorithm uses variance reduction, taking inspiration from methods such as Saga (Defazio et al.,, 2014) and Fedvarp (Jhunjhunwala et al.,, 2022). Our algorithm can be viewed as a relaxation of standard distributed stochastic gradient method, where our algorithm in each iteration shrinks the dissimilarities between the local models, instead of setting them to zero, i.e., averaging. Specifically, consider the following formulation of the Karula problem (1):
$\begin{array}[]{ll}\underset{\theta\in\mathcal{K}}{\mbox{minimize}}&f(\theta)=\frac{1}{n}\sum_{i=1}^{n}\alpha_{i}f_{i}(\theta_{i}),\quad\mathcal{K}=\{\theta\in{\mbox{\bf R}}^{d\times n}\colon\|\theta_{i}-\theta_{j}\|^{2}\leq tD_{ij}\}.\end{array}$
Suppose that each client $i$ may function as a stochastic first-order oracle $G_{i}$ that returns a stochastic gradient given an input model. With this formulation we can apply a stochastic projected gradient method, which is presented in Algorithm 1. In each training round of the algorithm, a subset of the clients return a stochastic gradient to the server. Subsequently, the server performs a variance reduced gradient step with respect to $f$, and projects the iterate onto $\mathcal{K}$.
The projection $\Pi_{\mathcal{K}}(\vartheta)$ of a collection of models $\vartheta\in{\mbox{\bf R}}^{d\times n}$ onto $\mathcal{K}$ does not have an analytical solution, so we need to employ an iterative method to compute it. Therefore, we define the $\delta$-inexact projection $\Pi_{\mathcal{K}}^{\delta}\colon\Theta^{n}\to\mathcal{K}$ that returns a feasible collection of models $\theta=\Pi_{\mathcal{K}}^{\delta}(\vartheta)$ which is at most $\delta$-suboptimal in the projection problem, meaning
@@
for the stepsize $\eta>0$. We give details on how to solve the projection computation in Appendix TODO.
@@ 4 Convergence analysis
In this section, we present an analysis of the convergence properties of our proposed algorithm for solving Problem (1), both in the general smooth case when the losses have Lipschitz continuous gradients, and in the smooth and strongly convex case. The analysis is inspired by the work of Reddi et al., (2016), which we have adapted for the FL setting, and to allow for inexact projections and stochastic gradients of each loss $f_{i}$.
In the smooth, possibly nonconvex case, we are interested in finding a local minimizer, but due to the problem being constrained, the ordinary stationarity condition $0=\nabla f(\theta)$ is not suitable. In lieu, the analogous condition for constrained optimization is $\theta$ being a zero of the gradient mapping
@@
To establish the convergence results, we make some standard assumptions. First, we assume smoothness of the losses.
@@
The weighted losses $\alpha_{i}f_{i}$ are $L$-smooth, meaning $\alpha_{i}\|\nabla f_{i}(\theta_{i})-\nabla f_{i}(\theta_{i}^{\prime})\|\leq L\|\theta_{i}-\theta_{i}^{\prime}\|$ for all $\theta_{i}$ and $\theta_{i}^{\prime}$.
Second, we will assume that the stochastic gradients have bounded mean squared error.
@@
The stochastic gradients returned by the oracles $G_{i}$ have bounded mean squared error, meaning ${\mathbbmss{E}}\|G_{i}(\theta_{i})-\alpha_{i}\nabla f_{i}(\theta_{i})\|^{2}\leq\sigma^{2}$ for every $\theta_{i}$.
Third, we will make use of strong convexity in one of the forthcoming convergence results.
Assumption 4.3.
The losses $f_{i}$ are $\gamma$-strongly convex, meaning $f_{i}(\theta_{i})\leq f_{i}(\theta_{i}^{\prime})+\langle\nabla f_{i}(\theta_{i}),\theta_{i}^{\prime}-\theta_{i}\rangle+(\gamma/2)\|\theta_{i}-\theta_{i}^{\prime}\|^{2}$.
We find under the two former assumptions that the proposed algorithm enjoys sublinear convergence to an $\epsilon$-stationary point, with the error $\epsilon$ arising due to the stochastic gradients and inexact projections.
@@
The assumption of feasible initialization is only made for the sake of exposition, and is trivially satisfied by initializing all clients to the same model.
Theorem 4.5.
Under Assumptions 4.1, 4.2 and 4.3, with feasible initialization, step size $\eta=3/(8L)$ and $12\leq s\leq n-$ 1,the iterates of Algorithm 1 satisfy for $k_{m}=mR\leq K$,
$\min_{k\in[k_{m},k_{m+1}]}{\mathbbmss{E}}[f(\theta^{k})-f(\theta^{\star})]\leq\frac{f(\theta^{0})-f(\theta^{\star})}{\rho^{m}}+\kappa\epsilon,$
where $\rho=\kappa R$ and $\epsilon=4\sigma^{2}(s+1)/s+2\delta$, and $\kappa=\gamma/L$.
@@
Note that in contrast to many other FL algorithms, our convergence analysis require no bounding of client heterogeneity for the convergence or other assumptions about the distributions, and instead relies only on standard machine learning assumptions. Moreover, the step size that is used to achieve the bound is comparatively genereous, and works well in practice when the smoothness constant $L$ is known or can be effectively approximated.
@@
While it is common to assume that the projections are exact, it is not particularly realistic. With our analysis, we explicitly account for error in the projection and show the contribution of the projection error to the neighborhood that the algorithm converges to. This informs practice, since we can choose the error $\delta$ to trade off the iteration cost with proximity to stationarity.
@@ 5 Numerical experiments
Here, we study the performance of Karula empirically and compare to some existing conventional and personalized FL strategies. First, in §5.1, we explore ridge regression on synthetic, heterogeneous data where the true model parameters are known. Then, in §5.2, we compare the different methods on a small-scale logistic regression task on heart disease data distributed accross four hospitals. Last, in §5.3, we evaluate the performance of the different methods on a federated handwritten digit classification task using neural network models.
@@
We compare Karula with five other FL strategies: local non-collaborative learning, federated stochastic gradient method, federated averaging, pFedMe and IFCA. For all methods, we weight the client loss functions by the number of samples.
@@ 5.1 Synthetic data
We begin by comparing the performance of the strategies on a ridge regression problem with $d=50$ features. The ridge penalization parameter is chosen as $\lambda=10^{-6}$, to ensure a unique solution. One third of the clients participate at every round and clients return full gradients.
@@
We generate a network of $n=30$ clients, where the ground-truth model parameters for a third of the clients are drawn from a normal distribution with mean $1$, another third from a normal distribution with mean $1.5$, and the remaining third from a normal distribution with mean $2$. We draw feature matrices with entries from three different normal distributions as well. Each client draws $N_{i}$ samples from their respective distribution, with $N_{i}$ being uniformly distributed between 10 and 100. We then set the underlying noise such that the global model performs slightly better than the completely local model.
Regularization parameter selection.
To set as level of a playing field as possible for the different FL strategies, we rely on 5-fold cross-validation to select the regularization parameters of pFedMe, IFCA and Karula. We do however acknowledge that it is not always feasible to do so in large-scale FL systems in practice.
@@ Results.
Table 1: Summary statistics of the performances of the different strategies on synthetic and heart disease data, with ±$\pm 2\times\mathrm{SE}$.
(a) Mean estimation error and test $R^{2}$ on synthetic data.
@@
(b) Test accuracy on heart disease data.
Strategy Test accuracy
Local 0.658
FedSGD 0.685
FedAvg 0.669
pFedMe 0.658
IFCA -
Karula 0.705
5.2 Federated heart disease data
We use the heart disease dataset provided by Janosi et al., (1988), which contains data on 940 patients spread out over four different hospitals in the US and in Europe. We preprocess the data as in (Ogier du Terrail et al.,, 2022), giving us 13 features and in total 740 samples, with a range of 46 to 303 patient samples per hospital.
@@ 5.3 Federated MNIST
We run experiments on a federated version of the MNIST dataset (Caldas et al.,, 2019). The task is classification, with ten different classes corresponding to digits $0,\dots,9$. Each client in the federated network owns data of handwritten digits corresponding to one individual writer. Here, heterogeneity in the network stems from idiosyncratic handwriting; some write small and others write large, some write “1” with serifs and others write without, etc.
@@ Results.
Figure 1:
@@ Appendix A Proofs of results
For simplicity, suppose without loss of generality that $\Theta={\mbox{\bf R}}^{p}$ for some natural number $p$.
@@
$f_{i}(\theta_{j}^{\star})-f_{i}(\theta_{i}^{\star})$ $\geq\frac{\mu}{2}\|\theta_{i}^{\star}-\theta_{j}^{\star}\|_{2}^{2},$
$f_{j}(\theta_{i}^{\star})-f_{j}(\theta_{j}^{\star})$ $\geq\frac{\mu}{2}\|\theta_{i}^{\star}-\theta_{j}^{\star}\|_{2}^{2}.$
@@
$\|\theta_{i}^{\star}-\theta_{j}^{\star}\|_{2}^{2}\leq\frac{1}{\mu}\left(f_{i}(\theta_{j}^{\star})-f_{j}(\theta_{j}^{\star})+f_{j}(\theta_{i}^{\star})-f_{i}(\theta_{i}^{\star})\right).$ (3)
By Kantorovich–Rubenstein Duality (Villani,, 2008), it holds for all $\theta$ that
@@
where Assumption 3.2 was used in the inequality, and where $\mathcal{L}_{1}$ is the space of $1$-Lipschitz functions. Combining this inequality with (3), we obtain the stated bound. ∎
A.2 Technical lemmatta
@@
Under Assumption 4.1 and 4.2, the iterates $\theta^{k}$, $\phi^{k}$ and $\nu^{k}$ of Algorithm 1 satsify
@@
$\Delta_{i}^{k}=\begin{cases}\pi_{i}(\nabla f_{i}(\theta_{i}^{k})-\nabla f_{i}(\phi_{i}^{k}))/s&\text{if }i\in S_{k},\\ 0&\text{otherwise},\end{cases}$
@@
In the the second inequality we used the bounded variance of the stochastic gradients (Assumption 4.2) and the Cauchy–Schwarz inequality coupled with Young’s inequality. In the third inequality we used $L$-smoothness of the weighted losses (Assumption 4.1). ∎
@@
holds for all $v\in\mathcal{K}$. Additionally,
@@
Moreover, since $1/(2\eta)\|v-(\theta-\eta\nu)\|^{2}$ is $1/(2\eta)$-strongly convex in $v$, for all $v\in\mathcal{K}$ the following inequality holds:
@@
After some rearranging, we obtain the inequality
@@
Using this coupled with (7), we obtain the stated inequality.
A.3 Proofs of Theorems
@@
where we used that $\mathbf{G}_{\eta}(\theta^{k})=(\theta^{k}-\vartheta^{k+1})/\alpha$. Now, we introduce the Lyapunov functions
$\nu^{k}={\mathbbmss{E}}\left[f(\theta^{k})+c_{k}\|\theta^{k}-\phi^{k}\|^{2}\right].$
@@
$=\nu^{k}+\left(L-\frac{1}{2\eta}\right){\mathbbmss{E}}\|\mathbf{G}_{\eta}(\theta^{k})\|^{2}+\frac{4(s+1)}{s}\sigma^{2}+2\delta.$
Summing over $k\in\{0,\dots,K-1\}$, and using the fact that $\nu^{k}=f(\theta^{K})$ due to $c_{K}=0$ and $V^{0}=f(\theta^{0})$ due to $\theta^{0}=\phi^{0}$, we arrive at the inequality

### v2

@@ 3.3 Algorithm
To solve the Karula problem (1), we introduce a projected gradient algorithm that supports partial client participation in each training round. To eliminate the error introduced by this partial participation, our algorithm uses variance reduction, taking inspiration from methods such as Saga (Defazio et al.,, 2014) and Fedvarp (Jhunjhunwala et al.,, 2022). Our approach can be viewed as a relaxation of standard distributed stochastic gradient method: rather than enforcing full mode agreement, each iteration shrinks model dissimilarities based on pairwise constraints. Specifically, we consider the following formulation of the Karula problem (1):
$\begin{array}[]{ll}\underset{\theta\in\mathcal{K}}{\mbox{minimize}}&f(\theta)=\frac{1}{n}\sum_{i=1}^{n}\alpha_{i}f_{i}(\theta_{i}),\quad\mathcal{K}=\{\theta\in\Theta^{n}\colon\|\theta_{i}-\theta_{j}\|^{2}\leq tD_{ij}\}.\end{array}$
Each client $i$ serves as a stochastic first-order oracle $G_{i}$ that returns a stochastic gradient when queried with a model $\theta_{i}$. Based on this formulation, we apply the stochastic projected gradient method outlined in Algorithm 1. In each training round, a subset of clients returns stochastic gradients to the server, which then performs a variance reduced gradient step with respect to $f$, and approximately projects the iterate onto the constraint set $\mathcal{K}$.
The projection $\Pi_{\mathcal{K}}(\vartheta)$ of a collection of models $\vartheta\in\Theta^{n}$ onto $\mathcal{K}$ does not admit an analytical solution, so we use an iterative method to approximate it. To this end, we define the $\delta$-inexact projection $\Pi_{\mathcal{K}}^{\delta}\colon\Theta^{n}\to\mathcal{K}$ that returns a feasible collection of models $\theta=\Pi_{\mathcal{K}}^{\delta}(\vartheta)$ which is at most $\delta$-suboptimal in the projection problem, meaning
@@
for the learning rate $\eta>0$. We describe how to perform the projection computation in the supplemental material.
@@ 4 Convergence analysis
In this section, we analyze the convergence properties of our proposed algorithm for solving Problem (1), both in the general smooth case when the losses have Lipschitz continuous gradients. The analysis is inspired by the work of Reddi et al., (2016), which we have adapted for the FL setting, and extended to allow for inexact projections and stochastic gradients of the individual client losses $f_{i}$.
Our goal in the smooth, possibly nonconvex case, is to find a local minimizer, but since the problem is constrained, the ordinary stationarity condition $0=\nabla f(\theta)$ is not appropriate. In lieu, the analogous condition for constrained optimization is that $\theta$ is a zero of the gradient mapping (cf. Nesterov,, 2018)
@@
To establish the convergence results, we make a few standard assumptions. First, we assume smoothness of the losses.
@@
The weighted losses $\alpha_{i}f_{i}$ are $L$-smooth, i.e.they are continuously differentiable with $L$-Lipschitz continuous gradients: $\alpha_{i}\|\nabla f_{i}(\theta_{i})-\nabla f_{i}(\theta_{i}^{\prime})\|\leq L\|\theta_{i}-\theta_{i}^{\prime}\|$ for all $\theta_{i}$ and $\theta_{i}^{\prime}$.
Second, we assume that the stochastic gradients have bounded mean squared error.
@@
The stochastic gradients returned by the oracles $G_{i}$ have bounded mean squared error: ${\mathbbmss{E}}\|G_{i}(\theta_{i})-\alpha_{i}\nabla f_{i}(\theta_{i})\|^{2}\leq\sigma^{2}$ for every $\theta_{i}$.
Under the two former assumptions, the proposed algorithm enjoys sublinear convergence to an $\epsilon$-stationary point where $\epsilon$ reflects the effect of stochastic gradients and inexact projections. Here, we assume feasible initialization only for the sake of exposition, which is trivially satisfied by assigning the same initial model to all clients.
@@
In contrast to many other FL algorithms, our convergence analysis does not require any bounds on client heterogeneity or assumptions about the underlying data distributions. Instead, it only relies on standard machine learning assumptions. Moreover, the step size that is used to achieve the bound is comparatively generous, and works well in practice provided that the smoothness constant $L$ is known or can be effectively estimated.
@@
While it is common to assume exact projections in theory, this is not particularly realistic. Our analysis explicitly accounts for projection errors and quantifies how these affect the size of the neighborhood that the algorithm converges to. This has practical implications, since it informs the choice of the error $\delta$ to trade off the iteration cost with proximity to stationarity.
@@ 5 Numerical experiments
We are now ready to study the performance of Karula empirically and compare it against several existing conventional and personalized FL strategies. First, in §5.1, we study ridge regression on synthetic, heterogeneous data where the true model parameters are known. Then, in §5.2, we evaluate the performance of the different methods on a federated handwritten digit classification task using neural network models. For simplicity, we use the Euclidean norm on the joint space of the features and labels for the dissimilarity parameter computation. The experiments in §5.1 and §5.2 were performed on an Intel Core i9 processor at 3.2 GHz, and on a cluster with Intel Xeon processors at 3.10 GHz, respectively.
@@
We compare Karula with three FL baselines: local non-collaborative learning, federated averaging with five local iterations, and IFCA. For all methods and experiments, we weight the client loss functions by the number of local samples.
@@ 5.1 Synthetic data
We begin by comparing the performance of the FL strategies on a ridge regression problem with $d=50$ features. The ridge penalty parameter is set to $\lambda=10^{-6}$ to ensure a unique solution. One third of the clients participate at every round, and clients return full gradients.
@@
We simulate a network of $n=30$ clients, where the ground-truth model parameters for a third of the clients are drawn from a normal distribution with mean $1$, another third from a normal distribution with mean $1.5$, and the final third from a normal distribution with mean $2$. Feature matrices are also drawn from distinct normal distributions across the clients. Each client draws $N_{i}$ samples from their assigned distribution, where $N_{i}$ is uniformly distributed between 10 and 100. The noise level is tuned so that the global model outperforms the completely local models.
Hyperparameter selection.
To set a level playing field for the comparison, we use 5-fold cross-validation to select the regularization parameters of IFCA and Karula. While this tuning approach may not always be practical in large-scale FL deployments, it provides a controlled benchmark setting for our experiments. The learning rate $\eta$ is set to $1/(10L)$, $1/(2L)$ and $3/(8L)$ for FedAvg, IFCA and Karula, respectively.
@@ Results.
Table 1 summarizes the results of the synthetic ridge regression experiments, reporting both estimation error and test $R^{2}$. Karula clearly outperforms the baselines, achieving the lowest estimation error and highest predictive performance. While FedAvg and IFCA improve substantially over fully local models, they fail to capture the underlying heterogeneity among clients. In particular, IFCA does not discover meaningful client clusters and effectively collapses to learning a single global model. To better understand this behavior, Figure 1 compares the true pairwise model dissimilarities with the dissimilarity parameters used by Karula and those estimated directly from local models. As shown, the dissimilarity structure inferred by Karula matches the ground truth closely, whereas direct estimation from local models fails to recover the correct relational structure. This result supports our argument from §3.2 that Wasserstein-based approximations provide more reliable measures of client similarity in low-sample regimes.
Table 1: Mean estimation error and test $R^{2}$ of the different strategies on synthetic ridge regression experiment, with ±$\pm 2\times\text{SE}$.
@@
(a) Ground truth
(b) Karula
(c) Local
Figure 1: Heatmaps showing the true model dissimilarities, the dissimilarity parameters calculated by Karula, and the model dissimilarities estimated via local learning, respectively.
@@ 5.2 Federated MNIST
Next, we run experiments on a federated version of the MNIST dataset (Caldas et al.,, 2019), where the task is to classify handwritten digits $\{0,\dots,9\}$. Each client holds data of handwritten digits corresponding to one individual writer. Here, heterogeneity stems from idiosyncratic handwriting: e.g., some write large, others small; some write “1” with serifs, others write without. We compare the performances of the different strategies on a network of $n=30$ clients, where half of them participate in every round. The clients return minibatch stochastic gradients with a batch size of 64.
Model architecture and personalization.
We use a two-layer perceptron model with ReLU activation and a hidden layer of 100 neurons. Following the personalization strategy in Arivazhagan et al., (2019) and the experimental setup of Ghosh et al., (2020), we apply personalization only to the second (output) layer of the model.
Hyperparameter selection.
For FedAvg and IFCA we use $\eta=10^{-4}$, whereas for the local model and Karula we use $\eta=2\times 10^{-4}$. For IFCA, we specify three clusters (following §6.3 in Ghosh et al.,, 2020). For Karula, the pairwise dissimilarity threshold is set to $t=20$.
@@ Results.
Figure 2 shows the training loss and test performance of the different strategies. The final average test accuracy is 91.3$\pm$0.75for Karula, 89.3$\pm$0.78for local learning, 84.4$\pm$2.39for FedAvg, and 83.9$\pm$2.60%, respectively. Notably, Karula not only yields better average accuracy, but also exhibits lower variance in the results, even though IFCA and FedAvg use smaller learning rates. To illustrate the interpretability of the dissimilarity parameters computed by Karula, Figure 3 visualizes the pairwise relationships among a subset of three clients. We clearly see how the learned dissimilarity structure reflects meaningful variations in the client data distributions, aligning with the underlying differences in handwriting.
(a) Training loss on federated MNIST.
(b) Test accuracy on federated MNIST.
Figure 2: Training and test performance with ±$\pm 2\times\text{SE}$ bands of FL strategies on classification task with 2NN model on Federated MNIST dataset with $n=30$ clients.
Figure 3: Example handwritten digits from three clients and the dissimilarity parameters $D_{ij}$ between them. The clients were chosen by randomly sampling one, and subsequently finding the nearest and most remote neighbor, respectively.
@@ 6 Discussion
We have introduced Karula, a personalized federated learning (PFL) strategy based on constrained optimization and data-driven measures of client dissimilarity. At the core of our approach is a surrogate for the 1-Wasserstein distance between client data distributions, which we compute efficiently via linear embeddings without any exchange of raw client data. This embedding-based dissimilarity offers a scalable and communication-efficient way to capture structural variation across clients. Under mild conditions on the loss functions, we proved that the distance between ideal personalized models can be bounded in terms of the Wasserstein distance between client distributions. This theoretical connection justifies constraining pairwise model dissimilarities as a means of improving generalization. To solve the resulting constrained problem, we proposed a projected stochastic gradient algorithm that supports partial client participation and incorporates variance reduction. We show that the method converges to an $\epsilon$-stationary point under standard assumptions, even for nonconvex losses. Empirical results on both synthetic and real-world federated datasets demonstrate the effectiveness of our approach. On the synthetic task, the learned dissimilarity parameters closely align with the true model distances, leading to superior personalized performance. On the federated MNIST benchmark, we could see that Karula can perform well on a real federated dataset, and that the algorithm converges quickly even when dealing with a nonconvex loss.
Limitations.
The model for communication constraints used in our analysis, although common in the federated learning literature, is not entirely realistic. It ignores the possibility of some clients communicating with the server significantly less often than others. Moreover, our accounting for inexact projections still assumes feasibility of the approximation, which is not always guaranteed in practice.
@@ Appendix A Proofs of results
$\frac{\mu}{2}\|\theta_{i}^{\star}-\theta_{j}^{\star}\|^{2}$ $\leq f_{i}(\theta_{j}^{\star})-f_{i}(\theta_{i}^{\star}),$
$\frac{\mu}{2}\|\theta_{i}^{\star}-\theta_{j}^{\star}\|^{2}$ $\leq f_{j}(\theta_{i}^{\star})-f_{j}(\theta_{j}^{\star})$
@@
$\|\theta_{i}^{\star}-\theta_{j}^{\star}\|^{2}\leq\frac{1}{\mu}\left(f_{i}(\theta_{j}^{\star})-f_{j}(\theta_{j}^{\star})+f_{j}(\theta_{i}^{\star})-f_{i}(\theta_{i}^{\star})\right).$ (3)
Define $\mathcal{L}_{1}$ as the space of real-valued $1$-Lipschitz functions on the data space $\mathcal{X}$. Then, by Kantorovich–Rubenstein Duality (Villani,, 2008, Theorem 5.10), it holds for all $\theta$ that
@@
where Assumption 3.2 was used in the inequality. Combining this inequality with (3), we obtain the stated bound. ∎
A.2 Technical lemmas
@@
Under Assumption 4.1 and 4.2, the iterates $\theta^{k}$, $\phi^{k}$ and $\nu^{k}$ of Algorithm 1 satisfy
@@
$\Delta_{i}^{k}=\begin{cases}\alpha_{i}(\nabla f_{i}(\theta_{i}^{k})-\nabla f_{i}(\phi_{i}^{k}))/s&\text{if }i\in S_{k},\\ 0&\text{otherwise},\end{cases}$
@@
In the second inequality we used the bounded variance of the stochastic gradients (Assumption 4.2) and the Cauchy–Schwarz inequality coupled with Young’s inequality. In the third inequality we used $L$-smoothness of the weighted losses (Assumption 4.1). ∎
@@
holds for all $v\in\mathcal{K}$. Moreover,
@@
Since $1/(2\eta)\|v-(\theta-\eta\nu)\|^{2}$ is $1/(2\eta)$-strongly convex in $v$, the following inequality holds for all $v\in\mathcal{K}$:
@@
Expanding squares and moving all terms to the right hand side, we obtain the inequality
@@
Using this coupled with (7), we obtain the stated inequality. ∎
A.3 Proofs of Theorem 4.3
@@
where we used that $\mathbf{G}_{\eta}(\theta^{k})=(\theta^{k}-\vartheta^{k+1})/\alpha$. Now, we introduce the Lyapunov function
$V^{k}={\mathbbmss{E}}\left[f(\theta^{k})+c_{k}\|\theta^{k}-\phi^{k}\|^{2}\right].$
@@
$=V^{k}+\left(L-\frac{1}{2\eta}\right){\mathbbmss{E}}\|\mathbf{G}_{\eta}(\theta^{k})\|^{2}+\frac{4(s+1)}{s}\sigma^{2}+2\delta.$
Summing over $k\in\{0,\dots,K-1\}$, and using the fact that $V^{K}=f(\theta^{K})$ due to $c_{K}=0$ and $V^{0}=f(\theta^{0})$ due to $\theta^{0}=\phi^{0}$, we arrive at the inequality